OpenStudy (anonymous):

Find the inverse of the following function:

OpenStudy (anonymous):

\[y=\log _{4}(2x-1)+3\]

OpenStudy (lochana):

Can you make an equation prior to x like y in your question

OpenStudy (jango_in_dtown):

replace x by y and y by x and then write x in terms of y

OpenStudy (anonymous):

So.... \[x=\log _{4}(2y-1)+3\]

OpenStudy (jango_in_dtown):

so we have \[x= \log_{4} (2y-1)+3 \]

OpenStudy (jango_in_dtown):

then \[(x-3)=\log_{4} (2y-1)\]

OpenStudy (lochana):

wow..sorry for the orientation. :D

OpenStudy (jango_in_dtown):

then \[4^{x-3}= 2y-1\]

OpenStudy (lochana):

@jango_IN_DTOWN you are correct. way to go!

OpenStudy (jango_in_dtown):

then \[y=(4^{x-3}+1)/2\]

OpenStudy (anonymous):

f^-1(x) = (4^x-3+1)/2

OpenStudy (jango_in_dtown):

now replace x by y and y by x again and write x as \[f ^-1 (x)\]

OpenStudy (jango_in_dtown):

it will give the inverse function